OpenStudy (anonymous):

The sales of a certain product after an initial release can be found by the equation given below, where s represents the total sales (in thousands) and t represents the time in weeks after release. Estimate the sales after 10 weeks. s=16sqrt(3t)+25

OpenStudy (anonymous):

calculator exercise put \(x=10\)

OpenStudy (mathstudent55):

Plug in 10 for t and calculate s.

OpenStudy (ddcamp):

Replace the t in the given equation with 10: 16×sqrt(3×10) + 25

OpenStudy (anonymous):

is that 3 times 10 or 3x10 @DDCamp

OpenStudy (anonymous):

it is 30

OpenStudy (anonymous):

would you subtract 25 on both sides?

OpenStudy (ddcamp):

No need to subtract, just plug it into your calculator and you'll get a decimal value for s.

OpenStudy (anonymous):

112.63?

OpenStudy (ddcamp):

Yup, but make sure you round correctly (the digit after the 3 is a 5) 112.635 → 112.64

OpenStudy (mathstudent55):

\(s = 16\sqrt{3t} + 25\) First, substitute 10 for t: \(s = 16\sqrt{3\times 10} + 25\) Do the multiplication inside the root. \(s = 16\sqrt{30} + 25\) Calculate teh root: \(s = 16 \times 5.477 + 25\) Multiply 16 and 5.477: \(s = 87.636 + 25\) Add \(s = 112.6 \) \( \text{Answer:} ~~$112,600 ~~\text{in sales} \)